OpenStudy (anonymous):

Solve |2y - 2| = y. A. {2/3, 2} B. {2, 3} C. {2, 5} D. {1/3, 2}

OpenStudy (anonymous):

2y-2=y and 2y-2=-y

OpenStudy (jackellyn):

2y -2 = y 2y-y = 2 y =2 2y-2 = -y 2y+ y =2 3y =2 y = 2/3

OpenStudy (anonymous):

thanks (: